PLease help with the following true/false and I am looking for examples if anyone has them..thank you so much
State whether the sentence is true or false. Give an example that
supports your answer.
(a) Zero cannot be the solution of an equation.
(b) A negative number cannot be the solution of an equation
(c) If an equation contains a negative number, then the solution of the
equation must be a negative number.

Answer by wuwei96815(245)

(a) False. What about quadratic equations? ax^2 + bx + c = 0

(b) False. 2x + y = -6, If x=-2 and y =6

(c) False. x - 3 = 9, If x = 12

I hope that is what you were looking for?